Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. What is the probability that the height of a child, picked at random, has a between 110 cm and 122 cm?


1
Expert's answer
2021-04-14T07:08:34-0400

Let "X" be a random variable denotes the height of the students.

Then "X" is normally distributed with mean "110" cm and standard deviation "6" cm: "X\sim N(110, 6^2)."


"P(110<X<122)=P(X<122)-P(X\leq110)"

"=P(Z<\dfrac{122-110}{6})-P(Z\leq\dfrac{110-110}{6})"

"=P(Z<2)-P(Z\leq0)"

"\approx0.97725-0.5=0.47725"



"P(110<X<122)=0.47725"